OpenStudy (anonymous):

On a baseball diamond, the distance from first base to third base is approximately 127 feet. How many inches is the distance from first base to third base?

OpenStudy (anonymous):

There are 12 inches for feet.

OpenStudy (anonymous):

So multiply 127 by 12 and you get your answer.

OpenStudy (anonymous):

It's 1,524.

OpenStudy (anonymous):

1'524 ^^

OpenStudy (anonymous):

127*12= 254+1270=1524

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

thanks for going over that with me

OpenStudy (anonymous):

Your welcome!